Calculate the molecular weight of 12.3g gas if 5.49L at 35°C hav3 pressure at 11.25 atm when R=0.082 L atm mol-1 K-1
1
Expert's answer
2020-06-23T02:16:55-0400

PV = nRT

n = PV / RT

T = 35 + 273.15 = 308.15 K

n = (11.25 x 5.49) / (0.082 x 308.15)

n = 2.444 mol

M = m/n

M = 12.3 g / 2.444 mol = 5.0 g/mol